# On the Self-awareness of Large Reasoning Models’ Capability Boundaries

### v2

@@ Abstract
Large Reasoning Models (LRMs) have shown impressive performance on complex reasoning tasks such as mathematics, yet they also display misbehaviors that expose their limitations. In particular, when faced with hard questions, LRMs often engage in unproductive reasoning until context limit, producing wrong answers while wasting substantial computation. This phenomenon reflects a fundamental issue: current answering paradigms overlook the relationship between questions and LRMs’ capability boundaries. In this paper, we investigate whether LRMs possess self-awareness of capability boundaries. We begin by an observation that LRMs may know what they cannot solve through expressed reasoning confidence. For black-box models, we find that reasoning expressions reveal boundary signals, with accelerated growing confidence trajectory for solvable problems but convergent uncertainty trajectory for unsolvable ones. For white-box models, we show that hidden states of the last input token encode boundary information, with solvable and unsolvable problems linearly separable even before reasoning begins. Building on these findings, we propose two simple yet effective optimization strategies: reasoning expression monitoring and hidden states monitoring. Experiments demonstrate that these boundary-aware strategies enable LRMs to avoid unproductive reasoning without sacrificing accuracy, significantly improving reliability and efficiency by cutting token usage up to – . We open-source at https://github.com/qingjiesjtu/CapBound.
@@ 1 Introduction
Large Reasoning Models (LRMs) have demonstrated remarkable capabilities on complex reasoning tasks such as mathematics (Guo et al., 2025; Jaech et al., 2024; Ahn et al., 2024). However, they also exhibit a range of misbehaviors that reveal the limitations of their capabilities (Kalai et al., 2025; Yao et al., 2025b; Sun et al., 2025; Zhang et al., 2024). In particular, when confronted with hard questions, LRMs may fall into repetitive looping until the context length is exhausted (Zhou et al., 2025; Chen et al., 2025b). Such a response paradigm not only fails to produce a correct answer but also wastes computational resources (Peng et al., 2025; Pu et al., 2025; Chen et al., 2024b), leaving users waiting a long time with no meaningful outcome.
This phenomenon highlights a fundamental issue: current answering paradigms of LRMs overlook the relationship between questions and their capability boundaries. For questions beyond their capability boundary, LRMs still struggle to produce a wrong answer, often engaging in unproductive long reasoning manifested as repetitive looping or error accumulation (Yao et al., 2025a; Mukherjee et al., 2025) (shown in Appendix E), which undermines both response efficiency and reliability. Therefore, it is crucial to identify questions beyond capability boundaries, and to explicitly acknowledge the inability to solve them. This importance has also been recognized by the broader AI community, as exemplified by the praise for GPT-5’s ability to appropriately say “I don’t know.”111https://x.com/koltregaskes/status/1957474061153436094
@@
Existing approaches to exploring capability boundaries are mainly conducted from an external perspective, for example, by constructing increasingly challenging benchmarks such as Humanity’s Last Exam (HLE) (Phan et al., 2025), or by designing real-time routers that quickly estimate question difficulty and assign harder questions to more capable models (OpenAI, 2025). However, little work has investigated capability boundaries from the model’s internal perspective, namely self-awareness of capability boundaries, which is particularly meaningful for understanding reasoning mechanism and for further improving reliability and efficiency.
Thus, this paper focuses on answering the question: “Do LRMs know their capability boundaries?” Figure 1 shows the overview of our work. We begin with the observation that LRMs often exhibit awareness of what they cannot solve (Section 3), and we provide evidence for such self-awareness for both black-box and white-box scenarios. For black-box models, we show that reasoning expressions contain clear signals of capability boundaries (Section 4). For white-box models, we find that hidden states encode information that also delineates these boundaries (Section 5). Taken together, these findings suggest that LRMs can assess, on the fly, whether a given question is beyond capability boundary, thereby enabling more efficient and reliable reasoning (Section 6). We summarize our contributions as follows.
We observe indications of LRMs’ self-awareness of capability boundaries by comparing their reasoning confidence, articulated in anthropomorphic tones, with the correctness of their final answers. Notably, LRMs know what they cannot solve.
From a black-box perspective, we reveal LRMs’ capability boundaries through reasoning trajectories. Remarkably, their reasoning shows an accelerated growing confidence trajectory for questions within the boundary, but a convergent uncertainty trajectory for questions beyond it.
@@
From a white-box perspective, we reveal LRMs’ capability boundaries through the hidden states of the last input token. Remarkably, solvable and unsolvable questions are almost linearly separable in the hidden space, even before the reasoning begins.
@@
To leverage capability boundaries during reasoning, we design two simple yet effective optimization strategies: Reasoning Expression Monitoring (black-box) and Hidden States Monitoring (white-box). Experiments show that these boundary-aware strategies substantially improve efficiency and reliability without sacrificing accuracy, reducing token usage up to – .
@@ 2 Related Work
From knowledge boundary to capability boundary. Recent studies have examined the knowledge boundaries of large language models (LLMs), focusing on the limits of factual knowledge encoded in parameters (Yin et al., 2024; Li et al., 2024). For instance, Wen et al. (2024) investigate LLMs’ perception of knowledge boundary through semi-open-ended question answering. Deng et al. (2025) unveil knowledge boundary of LLMs for trustworthy information access. Kale & Nadadur (2025) evaluate LLMs’ self-knowledge via consistency. These works primarily emphasize factual coverage, with retrieval-augmented generation (RAG) being a common solution to extend the accessible knowledge space (Gao et al., 2023; Lewis et al., 2020).
By contrast, the capability boundary concerns what reasoning tasks LLMs can reliably solve (Chen et al., 2025a; Xue et al., 2025; Chen et al., 2024a). This notion is broader yet rarely formalized. Existing efforts typically approximate it externally: either by constructing more challenging benchmarks such as Humanity’s Last Exam (HLE) (Phan et al., 2025), or by designing routing mechanisms such as the GPT-5 router that delegate harder questions to more capable models (OpenAI, 2025). In contrast, this paper investigates capability boundaries from the internal perspective of LLMs, focusing on their potential self-awareness.
@@
On the hidden states side, Wendler et al. (2024) probe multilingual Llama models and reveal English as the pivot language in the internal structure. Fan et al. (2024) demonstrate that certain intermediate layers already encode sufficient predictive features, implying redundancy in the full forward pass. Zhang et al. (2025) train linear probes on hidden states to verify intermediate answers, anticipating future correctness. Our work builds on both perspectives by analyzing reasoning expressions and hidden states to reveal self-awareness of capability boundaries.
@@
Efficient reasoning: from solvable questions to unsolvable questions. A range of recent works examine the phenomenon of overthinking (Sui et al., 2025). Chen et al. (2024b) show that LRMs often allocate excessive computation to easy questions, while Zhang et al. (2024) report that multi-turn self-correction can lead to unnecessary cost. To mitigate such inefficiency, Wang et al. (2025) demonstrate that suppressing frequent reflective tokens such as “Wait”, “Hmm”, or “Alternatively” improves reasoning efficiency without harming accuracy. Similarly, Yang et al. (2025a) introduce mechanisms to decide during reasoning whether further steps are unnecessary, allowing the model to terminate early and output the final answer directly.
@@
However, these works target the case where LRMs can solve a question but waste resources by reasoning too long. In contrast, our work addresses the complementary scenario: when a question is beyond the model’s capability boundary, LRMs still continue reasoning unproductively despite having no chance to get a correct answer.
3 Large Reasoning Models Know What They Cannot Solve
@@
We begin our exploration by evaluating commonly used LRMs including GPT-oss-20B (Agarwal et al., 2025), DeepSeek-R1-0528-Qwen3-8B (abbreviated as R1-Distill-Qwen3-8B) (Guo et al., 2025), DeepSeek-R1-Distill-Qwen-32B (abbreviated as R1-Distill-Qwen-32B) (Guo et al., 2025), and QwQ-32B (Qwen Team, 2024) on several standard mathematical benchmarks, including AIME’25 (MAA, 2025), AIME’24 (MAA, 2024), HMMT (HMMT, 2025), and AMC’23 (MAA, 2023) (see Appendix A for detailed descriptions of datasets and models).
@@
Inspired by works which monitors LRMs’ reasoning (Baker et al., 2025; He et al., 2025; Luo et al., 2024), we examine the reasoning processes (e.g., the segments wrapped within <think> in DeepSeek-like models) to look for signs of self-awareness regarding whether the input question can be correctly solved. As illustrated in Figure 2, we observe that such signals, expressed as anthropomorphic tones (Guo et al., 2025; Yang et al., 2025b), emerge throughout the reasoning process, and become especially evident in its final stages. Toward the end of reasoning, models often produce explicit statements that reflect their confidence level. Strikingly, these expressed beliefs tend to align with the correctness of their final answers. For example, when the reasoning contains more confident expressions such as “So, no mistake” or “I think it’s correct”, the final answers are typically correct. Conversely, when uncertain expressions like “My initial assumption is wrong” or “I’m not 100% sure” appear, the final answers are incorrect.
@@
This observation motivates us to quantify the alignment between expressed belief and answer correctness, which we formalize through a pair of metrics:
@@
Table 1 shows the results. We observe that LRMs’ expressed beliefs align with the correctness of their final answers. In particular, for questions where LRMs express more uncertainty during reasoning, the final answers are consistently incorrect (Cannot % = 100). This phenomenon is observed across multiple models, suggesting that LRMs are aware of what they cannot solve. Taken together, the reasoning expressions indicate a form of self-awareness of capability boundary.
4 Capability Boundary Revealed by Reasoning Expressions
Motivated by the observations in the previous section, we hypothesize that it is feasible to reveal LRMs’ capability boundaries through their reasoning expressions. Therefore, we systematically investigate the dynamics of confident and uncertain expressions throughout the reasoning process.
We first construct two sets of expressions, confident and uncertain expressions, that frequently appear in the reasoning process (the full list is provided in Appendix C). These anthropomorphic expressions are commonly observed and have been recognized in prior work as critical for reasoning (Yang et al., 2025b; Guo et al., 2025; Qian et al., 2025). Second, we divide the reasoning process into multiple stages (e.g., 50 stages) and compute the density of confident and uncertain expressions at each stage. Finally, concatenating these values yields the expression density trajectories, denoted as for confident expressions and for uncertain expressions. These trajectories capture how the density of each type of expression evolves over the course of reasoning.
We plot and in Figure 3, separately for cases with correct and wrong final answers. We observe that LRMs exhibit clear differences in their reasoning expression trajectories depending on final answer correctness. When the final answer is correct, the confident trajectory lies above the uncertain trajectory and shows an accelerating increasing trend. In contrast, when the final answer is incorrect, the uncertain trajectory dominates over the confident trajectory and tends to converge. Based on this distinction, the capability boundary can be defined as a linearly increasing trajectory (blue line in Figure 3). A concave trajectory indicates that the question lies within the LRM’s capability boundary, while a convex trajectory indicates that the question lies beyond it.
To formally characterize these patterns, we design two indicators as criteria for determining whether a question is within or beyond the capability boundary:
@@ 4 Capability Boundary Revealed by Reasoning Expressions
(2) -
Confidence Curvature (ConfCurv) computes the second derivative of the expression density trajectory to capture its convexity (Boyd & Vandenberghe, 2004). A concave trajectory corresponds to questions within the capability boundary, while a convex trajectory indicates questions beyond it. Also, we accumulate the sign of convexity up to a reasoning stage percentage and treat the question as beyond the capability boundary if the value exceeds a threshold :
Figure 4 shows that the two proposed indicators can reliably distinguish unsolvable from solvable questions, achieving high accuracy across different LRMs throughout the reasoning stage. Between the two indicators, ConfDiff is more reliable as it achieves higher accuracy and exhibits smaller fluctuations during the reasoning process. Moreover, the high accuracy emerges early, e.g., as early as 2% for GPT-oss-20B. Interestingly, the gradually declining yellow curve of this model suggests that boundary signals are more salient at the beginning than at the end of reasoning. This is likely due to repetitive looping and error accumulation in later stages (Yao et al., 2025a; Mukherjee et al., 2025). Therefore, we demonstrate that the capability boundary can be revealed through reasoning expressions, even at an early stage.
5 Capability Boundary Revealed by Hidden States
Following the observation that capability boundaries can be revealed at an early stage of reasoning, it is worth examining whether such signals may emerge even earlier, potentially before the model begins reasoning. Therefore, we investigate whether capability boundaries are already encoded in the hidden states of input questions.
Since hidden states are high-dimensional (e.g., 4096), limited samples may lead to sparsity issues that obscure statistical patterns (Vishwakarma, 2024; Köppen, 2000). To mitigate this, we first expand our dataset by including GSM8K (Cobbe et al., 2021) and Humanity’s Last Exam (HLE) (Phan et al., 2025). These two benchmarks provide complementary levels of difficulty (see Appendix A for full accuracy results): GSM8K is relatively easy, with an accuracy exceeding 90%; HLE is extremely difficult, with an accuracy below 10%. This contrast helps amplify the distinction between solvable and unsolvable questions, making capability boundaries more salient.
Next, we extract hidden states from the prefilling stage as it processes the input question (Yuan et al., 2024; Pope et al., 2023), using the hook method (Vig, 2019). Specifically, we focus on the hidden state of the last input token, which encapsulates the semantic content of the entire question (Radford et al., 2019). Finally, we apply simple linear classifiers to these hidden states, including Linear Discriminant Analysis (LDA) and Logistic Regression (LR) (Hastie, 2009). Using final answer correctness as labels, we test whether solvable and unsolvable questions can be linearly separated, thereby revealing the presence of capability boundaries in hidden states.

### v3

# Stop Before You Fail: Operational Capability Boundaries for Mitigating Unproductive Reasoning in Large Reasoning Models

@@ Abstract
Current answering paradigms for Large Reasoning Models (LRMs) often fail to account for the fact that some questions may lie beyond the model’s operational capability boundary, leading to long but unproductive reasoning. In this paper, we study whether LRMs expose early signals predictive of such cases, and whether these signals can be used to mitigate unproductive reasoning. In black-box settings, we find that reasoning expressions contain failure-predictive signals. In white-box settings, we show that the hidden states of the last input token contain information that is predictive of whether a question will not be solved correctly under our evaluation setup. Building on these observations, we propose two test-time monitoring strategies: reasoning expression monitoring and hidden states monitoring, that reduce token usage by 62.7–93.6%, substantially improving efficiency and reliability while largely preserving accuracy.111We open-source our code.
Stop Before You Fail: Operational Capability Boundaries for Mitigating Unproductive Reasoning in Large Reasoning Models
Qingjie Zhang1, Yujia Fu1, Yang Wang2, Liu Yan2,
Tao Wei2, Ke Xu1, Minlie Huang1, and Han Qiu1††thanks: The corresponding author
1Tsinghua University, 2Ant Group
Emails:{qj-zhang24@mails., qiuhan@}tsinghua.edu.cn
@@ 1 Introduction
Large Reasoning Models (LRMs) have demonstrated remarkable capabilities on complex reasoning tasks such as mathematics (Guo et al., 2025; Jaech et al., 2024; Ahn et al., 2024). However, when confronted with hard questions, they frequently exhibit unproductive reasoning, manifested as repetitive looping or error accumulation until context limits are exhausted (Zhou et al., 2025; Chen et al., 2025b; Yao et al., 2025; Mukherjee et al., 2025) (as shown in Appendix K).
This phenomenon suggests an important limitation of current answering paradigms: they often fail to account for the fact that some questions are likely to fall beyond the model’s operational capability boundary under a fixed evaluation setup, leading to long yet ultimately unproductive reasoning traces. If such cases can be identified early, reasoning could become more reliable and efficient: instead of continuing low-value deliberation, the model may abstain from full reasoning and provide a concise partial approach.
@@
Existing approaches to studying reasoning limits have primarily taken an external perspective, for example by constructing increasingly challenging benchmarks such as Humanity’s Last Exam (HLE) (Phan et al., 2025), or by designing real-time routers that estimate question difficulty and route harder questions to more capable models (OpenAI, 2025). In contrast, relatively little work has examined whether such limits are reflected in the model’s internal signals, despite the importance of this question for both understanding reasoning behavior and further improving reliability and efficiency at test time.
Accordingly, this paper studies whether LRMs expose early signals predictive of eventual failure, and whether such signals can be used to mitigate unproductive reasoning. We use the term capability boundary in an operational sense: under a fixed evaluation setup, it refers to the empirical separation between questions the model eventually solves and those it does not. Figure 1 provides an overview of our approach. From both black-box and white-box perspectives, we identify signals that help distinguish unsolvable from solvable questions. Leveraging these signals, we enable LRMs to assess, on the fly, whether a question is likely to remain unsolved under the current setup, thereby improving the efficiency and reliability of reasoning. We summarize our contributions as:
We provide empirical evidence that LRMs’ reasoning expressions are associated with the correctness of final answers, revealing language patterns that are predictive of eventual failure.
From a black-box perspective, we characterize failure-predictive reasoning trajectories. We identify a diverging pattern: an accelerated growing confidence trajectory for solved questions, but a convergent uncertainty trajectory for unsolved questions.
@@
From a white-box perspective, we show that the hidden states of the last input token contain information predictive of whether a question will eventually be solved correctly. In particular, solvable and unsolvable questions are close to linearly separable in the hidden space, even before reasoning begins.
@@
We design two test-time monitoring strategies: reasoning expression monitoring (black-box) and hidden states monitoring (white-box). Experiments show that these strategies substantially improve efficiency and reliability while largely preserving accuracy, reducing token usage by –.
@@ 2 Related Work
From knowledge boundary to capability boundary. Recent studies have examined the knowledge boundaries of Large Language Models (LLMs), focusing on the limits of factual knowledge encoded in parameters (Yin et al., 2024; Li et al., 2024). For instance, Wen et al. (2024) investigate LLMs’ perception of knowledge boundary through semi-open-ended question answering. Deng et al. (2025) unveil knowledge boundary of LLMs for trustworthy information access. These works primarily emphasize factual coverage, with retrieval-augmented generation (RAG) being a common solution to extend the accessible knowledge space (Gao et al., 2023; Lewis et al., 2020).
By contrast, the capability boundary concerns what reasoning tasks LLMs can reliably solve (Chen et al., 2025a; Xue et al., 2025; Chen et al., 2024a). This notion is broader yet rarely formalized. Existing efforts typically approximate it externally: either by constructing more challenging benchmarks such as Humanity’s Last Exam (HLE) (Phan et al., 2025), or by designing routing mechanisms such as GPT-5 router that delegate harder questions to more capable models (OpenAI, 2025). In contrast, we examine whether capability boundaries are reflected in models’ internal signals.
@@
On the hidden states side, Wendler et al. (2024) probe multilingual Llama models and reveal English as the pivot language in the internal structure. Fan et al. (2024) demonstrate that certain intermediate layers already encode sufficient predictive features, implying redundancy in the full forward pass. Zhang et al. (2025) train linear probes on hidden states to verify intermediate answers, anticipating future correctness. Building on both perspectives, our work analyzes reasoning expressions and hidden states to study signals related to the operational capability boundary.
@@
Efficient reasoning: from solvable questions to unsolvable questions. A range of recent works examine the phenomenon of overthinking (Sui et al., 2025; Hou et al., 2025). Chen et al. (2024b) show that LRMs often allocate excessive computation to easy questions, while Zhang et al. (2024) report that multi-turn self-correction can lead to unnecessary cost. To mitigate such inefficiency, Han et al. (2025) compresses unnecessary reasoning by including a reasonable token budget in the prompt. Wang et al. (2025) demonstrate that suppressing frequent reflective tokens such as “Wait”, “Hmm”, or “Alternatively” improves reasoning efficiency without harming accuracy. Yang et al. (2025a); Fu et al. (2025, 2024) introduce mechanisms to decide during reasoning whether further steps are unnecessary, allowing the model to terminate early and output the final answer directly.
@@
However, these works focus on cases where LRMs can solve a question but still waste computation by reasoning for too long. By contrast, our work addresses a complementary and less-studied scenario: questions that are likely to fall beyond the model’s operational capability boundary, where LRMs may nevertheless continue reasoning unproductively despite being unlikely to arrive at a correct answer under the current setup.
3 LRMs Expose Signals of Failure
@@
We begin our exploration by evaluating GPT-oss-20B (Agarwal et al., 2025), DeepSeek-R1-0528-Qwen3-8B (abbreviated as R1-Distill-Qwen3-8B) (Guo et al., 2025), DeepSeek-R1-Distill-Qwen-32B (abbreviated as R1-Distill-Qwen-32B) (Guo et al., 2025), and QwQ-32B (Qwen Team, 2024) on several standard mathematical benchmarks, including AIME’25 (MAA, 2025), AIME’24 (MAA, 2024), HMMT (HMMT, 2025), and AMC’23 (MAA, 2023) (see Appendix A for detailed descriptions of datasets and models).
@@
Inspired by works which monitors LRMs’ reasoning (Baker et al., 2025; He et al., 2025; Luo et al., 2024), we examine the reasoning processes (e.g., the segments wrapped within <think> in DeepSeek-like models) to look for internalized assessments regarding whether the input question can be correctly solved. As illustrated in Figure 2, we observe that such signals, expressed as anthropomorphic tones (Guo et al., 2025; Yang et al., 2025b), emerge throughout the reasoning process, and become especially evident in its final stages. Toward the end of reasoning, models often produce explicit statements that reflect their confidence level. Strikingly, these expressed beliefs tend to align with the correctness of their final answers. For example, when the reasoning contains more confident expressions such as “So, no mistake” or “I think it’s correct”, the final answers are typically correct. Conversely, when uncertain expressions like “My initial assumption is wrong” or “I’m not 100% sure” appear, the final answers are incorrect.
@@
This observation motivates us to quantify the alignment between expressed belief and answer correctness, which we formalize through:
@@
Table 1 shows the results. We observe that LRMs’ expressed beliefs align with the correctness of their final answers. In particular, for questions where LRMs express more uncertainty during reasoning, the final answers are consistently incorrect (Cannot % = 100). This phenomenon is observed across multiple models, suggesting that LRMs expose signals of failure.
4 Capability Boundary in Expressions
Motivated by the observations in the previous section, we hypothesize that LRMs’ reasoning expressions may contain signals related to the operational capability boundary. We therefore systematically investigate the dynamics of confident and uncertain expressions throughout the reasoning process.
To ensure that the analysis is not tied to any particular LRM’s reasoning style, we first extract confident and uncertain expressions separately for each model, based on its characteristic reasoning patterns (see the detailed extraction pipeline in Appendix C). Such anthropomorphic expressions are commonly observed in LRM reasoning traces and have been discussed in prior work as informative features of the reasoning process (Yang et al., 2025b; Guo et al., 2025; Qian et al., 2025).
Then, we uniformly divide the reasoning process into multiple stages (e.g., 50 stages) and compute the density of confident and uncertain expressions at each stage. Finally, concatenating these values yields the expression density trajectories, denoted as for confident expressions and for uncertain expressions. These trajectories capture how the density of each type of expression evolves over the course of reasoning.
We plot and in Figure 4, separately for cases with correct and wrong final answers. We observe that LRMs exhibit clear differences in their reasoning expression trajectories depending on final answer correctness. When the final answer is correct, the confident trajectory lies above the uncertain trajectory and shows an accelerating increasing trend. In contrast, when the final answer is incorrect, the uncertain trajectory dominates over the confident trajectory and tends to converge. This contrast suggests that reasoning expression trajectories contain useful signals (blue line in Figure 4) for distinguishing questions that are eventually solved from those that remain unsolved under the current setup. In particular, the dominant trajectory tends to be more concave for solvable questions and more convex for unsolvable ones.
To formalize these patterns, we design two indicators for determining whether a question is likely to lie within or beyond the capability boundary.
@@ 4 Capability Boundary in Expressions
Confidence Curvature (ConfCurv) computes the second derivative of the expression density trajectory to capture its convexity (Boyd and Vandenberghe, 2004). A concave trajectory corresponds to questions within the capability boundary, while a convex trajectory indicates questions beyond it. Also, we accumulate the sign of convexity up to a reasoning stage percentage and treat the question as beyond the capability boundary if the value exceeds a threshold :
Figure 3 shows that the two proposed indicators can reliably distinguish unsolvable from solvable questions, achieving high accuracy across different LRMs throughout the reasoning stage. Between the two indicators, ConfDiff is more reliable as it achieves higher accuracy and exhibits smaller fluctuations during the reasoning process. Moreover, strong predictive accuracy emerges early, e.g., as early as 2% for GPT-oss-20B. Interestingly, the gradually declining yellow curve of this model suggests that these signals may be more salient at the beginning than at the end of reasoning. One possible explanation is that repetitive looping and error accumulation in later stages make the trajectories noisier (Yao et al., 2025; Mukherjee et al., 2025). Overall, these results show that reasoning expressions contain early signals related to the operational capability boundary and predictive of eventual failure.
5 Capability Boundary in Hidden States
Following the observation that reasoning expressions expose early signals related to the operational capability boundary, it is natural to ask whether similar signals may emerge even earlier, potentially before the model begins reasoning. We therefore investigate whether hidden states of the input question already contain information predictive of eventual success or failure.